Use Kami in completing this activity.

The Sun has a mass of 2.0 x 10 to the power of 30 kg. Its

radius is 7.0 x 10 to the power of 8 meters. The gravitational

force between the sun and a section of the sun's surface is

470 N. What is the mass of this section?



1
Expert's answer
2021-12-23T10:49:41-0500

The mass of this section can be found from Newton's second law:


"F=G\frac{Mm}{R^2},\\\space\\
m=\frac{FR^2}{MG}=1.73\text{ kg}."